Growth​ Company's current share price is $20.30 and it is expected to pay a $0.85 dividend per share next year. After​ that, the​ firm's dividends are expected to grow at a rate of 4.2% per year.

a. What is an estimate of Growth​ Company's cost of​ equity?

b. Growth Company also has preferred stock outstanding that pays a $2.05 per share fixed dividend. If this stock is currently priced at $28.15​, what is Growth​ Company's cost of preferred​ stock?

c. Growth Company has existing debt issued three years ago with a coupon rate of 5.8%. The firm just issued new debt at par with a coupon rate of 6.4%. What is Growth​ Company's cost of​ debt?

d. Growth Company has 4.7 million common shares outstanding and 1.4 million preferred shares​outstanding, and its equity has a total book value of $50.3 million. Its liabilities have a market value of $19.6 million. If Growth​ Company's common and preferred shares are priced as in parts (a​) and (b​), what is the market value of Growth​ Company's assets?

e. Growth Company faces a 40% tax rate. Given the information in parts (a​) through (d​), and your answers to those​ problems, what is Growth​ Company's WACC?

​Note: Assume that the firm will always be able to utilize its full interest tax shield.
Answer #1

a). kE = [D1 / P0] + g = [$0.85 / $20.30] + 0.042 = 0.0419 + 0.042 = 0.0839, or 8.39%

b). kP = Annual Dividend / Current Market Price = $2.05 / $28.15 = 0.0728, or 7.28%

c). As the latest bond issue has a kD of 6.4%(Issued at par). So, the company's kD is 6.4%

d). Market Value of Debt = $19.6 million

Market Value of Common Equity = $20.30 x 4.7 million = $95.41 million

Market Value of Preferred Equity = $28.15 x 1.4 million = $39.41 million

Total Market Value of Assets = 19.6 + 95.41 + 39.41 = $154.42 million

WACC = [wD x kD x (1 - t)] + [wP x kP] + [wE x kE]

= [(19.6/154.42) x 6.4% x (1 - 0.40)] + [(39.41/154.42) x 7.28%] + [(95.41/154.42) x 8.39%]

= 0.49% + 1.86% + 5.18% = 7.53%